4 a) joy, using, implicit differentakel 6) If y = (4x+5) - ² + ln (4x+6) - ex-8 then sind go (derivative of
Answer #1

** (356 + 7) = 64 da ( 324 +7)* -de6y) 2(324+7)(34+3%Y)=640 gxuy?+ 9x{y+214+2/xy= 34 y( 9x?+21X—3)=- grey #217 y=-9xy-214